OpenStudy (hypersniper445):

Please help I will fan and medal! Complete the standard form of the equation –5y = 10 – 2x. ???x + ???y = 10

OpenStudy (hypersniper445):

@daggiemib help

OpenStudy (daggiemib):

add 2x on both sides

OpenStudy (hypersniper445):

2x + 5y = 10?

OpenStudy (daggiemib):

2x-5y=10

OpenStudy (hypersniper445):

i cant type that in

OpenStudy (hypersniper445):

for some reason it has to be a plus

OpenStudy (hypersniper445):

do i just make the 5 negative?

OpenStudy (daggiemib):

well the sign shouldnt be positive...

OpenStudy (hypersniper445):

i can send a screenshot i cant change it

OpenStudy (daggiemib):

ok

OpenStudy (hypersniper445):

it is the same for all of the answers. it has a plus but everyone comes up with a minus instead

OpenStudy (daggiemib):

can u put in (-5)

OpenStudy (hypersniper445):

yes

OpenStudy (daggiemib):

2x+(-5)y=10

OpenStudy (hypersniper445):

okay thank you. could you help with a few more?

OpenStudy (daggiemib):

sure

OpenStudy (hypersniper445):

cool one second

OpenStudy (daggiemib):

ok give me a second to write it out

OpenStudy (hypersniper445):

my guess is C if that helps

OpenStudy (daggiemib):

i got c

OpenStudy (hypersniper445):

okay good. i think there are 3 more. is that okay?

OpenStudy (daggiemib):

no prob

OpenStudy (daggiemib):

i got -5.3

OpenStudy (hypersniper445):

hmm I got 2... could you explain please?

OpenStudy (hypersniper445):

its okay, i think i got it

OpenStudy (hypersniper445):

okay i see

OpenStudy (daggiemib):

do u want a pic of the other one?

OpenStudy (hypersniper445):

no its okay. ill give you the next one.

OpenStudy (hypersniper445):

I guess line D

OpenStudy (daggiemib):

b y=-2/5x+2

OpenStudy (daggiemib):

remember slope is rise over run

OpenStudy (hypersniper445):

oh yeah okay i get it

OpenStudy (hypersniper445):

last one

OpenStudy (daggiemib):

idk what collinear means give me a moment

OpenStudy (hypersniper445):

k

OpenStudy (daggiemib):

col·lin·e·ar kəˈlinēər,kä-/Submit adjectiveGEOMETRY (of points) lying in the same straight line.

OpenStudy (hypersniper445):

D right?

OpenStudy (daggiemib):

so given that i am guess its (0,0)

OpenStudy (hypersniper445):

oh sorry i meant line d so yeah a

OpenStudy (hypersniper445):

wait i just confused myself

OpenStudy (daggiemib):

its in-between line c and b so you understand?

OpenStudy (hypersniper445):

okay yeah i see

OpenStudy (daggiemib):

point not line lol sorry

OpenStudy (hypersniper445):

okay let me see what i got. anything you need to double check before i do so?

OpenStudy (hypersniper445):

*we* sorry

OpenStudy (daggiemib):

no im pretty sure their correct

OpenStudy (hypersniper445):

same

OpenStudy (hypersniper445):

Yay! I got 100% Thank you so much I failed that lesson 3 times because no one in our family knew how to do it lol thanks you are awesome :)

OpenStudy (daggiemib):

aw yay! glad i could help you out!:) tag me if you need anymore help

OpenStudy (hypersniper445):

will do! Have a nice day